OpenStudy (thephysicsman):

I was looking at functions and their inverses: is this always true: A function and it's inverse are a reflection over the line y = x

jimthompson5910 (jim_thompson5910):

yes because x and y swap places

OpenStudy (thephysicsman):

so like what if you take a function and find it's inverse and they are not a reflection around the line y = x. that does not make them inverses right?

jimthompson5910 (jim_thompson5910):

then you made a mistake somewhere in finding the inverse

OpenStudy (thephysicsman):

hmm

OpenStudy (thephysicsman):

This was a problem well kind of finding the inverse for \[y = Ae^{-kx}\]

OpenStudy (thephysicsman):

I just made this up lol

OpenStudy (thephysicsman):

\[y = 2e^{-2x}\]

jimthompson5910 (jim_thompson5910):

Swap x and y. Then solve for y \[\Large y = Ae^{-kx}\] \[\Large x = Ae^{-ky}\] \[\Large \frac{x}{A} = e^{-ky}\] \[\Large e^{-ky} = \frac{x}{A}\] \[\Large -ky = \ln\left(\frac{x}{A}\right)\] \[\Large y = -\frac{1}{k}*\ln\left(\frac{x}{A}\right)\]

OpenStudy (thephysicsman):

you are the man

OpenStudy (thephysicsman):

thanks I'm gonna graph this to see how they are related.

jimthompson5910 (jim_thompson5910):

it's probably not 100% obvious, but the orange and purple curves are mirror images of each other https://www.desmos.com/calculator/igmdna30pd one way to check is to pick a point on the purple curve and reflect it over to find a corresponding point on the orange curve. Do this with a bunch of points to get a better sense of the reflection going on

OpenStudy (thephysicsman):

I'm going to do that now

jimthompson5910 (jim_thompson5910):

For instance (0,1) on the purple curve reflects over to (1,0) on the orange curve. Notice how x and y swapped places.

OpenStudy (thephysicsman):

Yeah what I did was I plotted (2,0.0185) and then (0.0185,2) and it was on the inverse. so it seems that the x and y points are swapped.

OpenStudy (thephysicsman):

so it seems like every point on f(x) (x,y) becomes (y,x) on f^-1(x)

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

because of the coordinate swap, the domain and range also switch too